OpenStudy (swag):

Given the points (2,1) and (3,4) A) What is the slope B) Write in point slope form using the slope you found in part A and one of the ordered pairs given. It doesn’t matter which ordered pair you use. The point slope formula is (y – y1) = m(x – x1). C) Simplify your answer in part B and write the equation in slope-intercept form. Show work please so i understand.

OpenStudy (anonymous):

Hey!:) Do you have any idea how to work this problem or are you completely lost?

OpenStudy (swag):

COMPLETELY, I was hoping it could be completely done for me so i could understand. This is a question on my study guide so it is not worth anything but I want to know how to do it.

OpenStudy (anonymous):

Cool, I can help you out. :)

OpenStudy (swag):

Ok thanks :)

OpenStudy (anonymous):

It'll take me a couple of minutes to type it out, but here goes:

OpenStudy (swag):

Np

OpenStudy (anonymous):

I'm going to show you part a first, so you don't get confused.

OpenStudy (swag):

Ok

OpenStudy (anonymous):

Given the points (2,1) and (3,4) A) What is the slope In order to find the slope, you need to follow this formula: slope (or, m) = (y2 - y1)/(x2 - x1) The y values in this problem are 1 and 4. You can pick either one of these values to represent y1. I'm going to use 1. Whichever value you don't use is going to be y2. (4 - 1)/(x2 - x1) It's not quite the same with the x values, which are 2 and 3. Once you have picked a y value to represent y1, you have to use the matching x. You can't use onne from the other coordinate, or you will screw up the answer. So in this case, x1 is 2. (4 - 1)/(3 - 2) Now solve the little equations before you divide: 4 - 1 = 3 and 3 - 2 = 1 That means that your answer looks like this: 3/1 or just 3. That is your slope. Do you get that?

OpenStudy (swag):

Yes so far

OpenStudy (anonymous):

Okay, cool. So if you feel like you are comfortable with finding the slope of a line, we can move onto part b?

OpenStudy (swag):

Yes

OpenStudy (anonymous):

Cool. Let me finish typing this up. Let me know if you need me to get a move on it, though, okay? :)

OpenStudy (swag):

ok

OpenStudy (anonymous):

m= y2-y1/x2-x1 m=1-4/2-3 m= -3/-1=3 y-4=3 (x-3) y-4= 3x-9 y= 3x-5

OpenStudy (anonymous):

So far, we have (2,1) and (3,4) as our coordinates, and a slope of 3. B) Write in point slope form using the slope you found in part A and one of the ordered pairs given. It doesn’t matter which ordered pair you use. The point slope formula is (y – y1) = m(x – x1). So for point-slope form, you basically just have to plug in values for the variables in the point slope formula: (y – y1) = m(x – x1) You can choose any of the two coordinates they gave you. I'll pick (2,1) for this example. Remember that "m" is the variable that means slope. (Seems kind of silly to me, where the heck is "m" in the word slope?!) Our slope is 3. Okay, so this is what it will look like: y - 1 = 3(x - 2) Then we solve. y - 1 = 3x - 6 (add +1 to both sides) y = 3x - 5 And that's your answer. Do you get that concept, now?

OpenStudy (swag):

Ok I get it but just to be clear can you break them down a little more clear and label A B AND C

OpenStudy (anonymous):

Yes. Here, let me break it down for you: PART B: B) "Write in point slope form using the slope you found in part A and one of the ordered pairs given. It doesn’t matter which ordered pair you use. The point slope formula is (y – y1) = m(x – x1)." So for point-slope form, you basically just have to plug in values for the variables in the point slope formula: (y – y1) = m(x – x1) You can choose any of the two coordinates they gave you. I'll pick (2,1) for this example. Remember that "m" is the variable that means slope. (Seems kind of silly to me, where the heck is "m" in the word slope?!) Our slope is 3. Okay, so this is what it will look like in slope intercept (answer to part B): y - 1 = 3(x - 2) PART C: Then we solve the point-slope to get the slope intercept. y - 1 = 3x - 6 (add +1 to both sides) y = 3x - 5 ^^ANSWER to part C So do you see that more clearly now?

OpenStudy (anonymous):

Do you still need part A? I did Part A by itself above; and you said you got it...

OpenStudy (swag):

Yes let me review it

OpenStudy (anonymous):

Sure! :) Just let me know

OpenStudy (swag):

This is perfect thank you so much

OpenStudy (anonymous):

You are welcome! I'm glad I could help! :D